OpenStudy (anonymous):

how do you explain how i got the awnser to x to the power of 4 over x to the power of 2

OpenStudy (anonymous):

\[\frac{ X ^{4} }{ X ^{2}}=X ^{4-2}=X ^{2}\]

OpenStudy (triciaal):

when you divide numbers with the same base subtract the powers x^4/x^2 = x^(4-2) = x^2 when you divide you are cancelling like factors x*x*x*x / x*x = x*x